It takes a plane 5 hours to travel 3000 miles with a tailwind, but it takes the same plane 6 hours to travel the same distance with a headwind. What is the speed of the plane in still air and the speed of the wind?
Answer by stanbon(75887)

It takes a plane 5 hours to travel 3000 miles with a tailwind, but it takes the same plane 6 hours to travel the same distance with a headwind. What is the speed of the plane in still air and the speed of the wind?
-------
Let plane speed be p and wind speed be w.
--------
With the wind DATA:
distance = 3000 mi ; time=5 hrs ; rate= d/t= 600 mph
---------
Against the wind DATA:
distance = 3000 mi ; time= 6 hrs; rae = d/t = 500 mph
---------
EQUATION:
p+w=600 mph
p-w=500 mph
Add the wo quations to get:
2p=1100
p=550 mph (plane's speed in still air)
Substitute to solve for w, as follows:
550+w=600
w=50 mph (wind's speed)
Cheers,
Stan H.